OpenStudy (anonymous):

In a right triangle, if one acute angle has measure 6x-10 and the other has measure 2x-4 , find the larger of these two angles.

OpenStudy (anonymous):

We can set these two angles = 90degrees, since we know the third angle is 90degrees and triangles sum to 180degrees. So: 6x-10 + 2x-4= 90 Now we can add like terms and solve for x: 8x-14 = 90 8x= 104 x= 13 Now we can sub x back into the given equations and see which angle is the largest: 6(13) - 10 = 68degrees and 2(13) - 4 = 22degrees